

********************************************************************
UARS WEEKLY STATUS REPORT - 24 MAR 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG worked with the Flight Operations Team to
develop daily tape recorder operation schedules
so that the maximum amount of northern hemisphere
data will be acquired by MLS data during the operation
of MLS from March 26 through March 30, 2000.
Routine processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
STATUS:
- The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
5 days HAL (4.82) PROD MODE (L1-L3) (23R-23C=0L)
- The CDHF completed processing the SOLVE Mission days (3083 - 3105)
requested by HALOE.
- Waiting further instructions concerning WINDII V5.11A Software.
All requested days are complete.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, SUSIM,
and Solstice-(HOORAY).
- SOLSTICE V4.62 (JOB 1 and JOB 2) was successful on the Rehost System.
Files and Sizes were compared to files that were created on the old
system. SOLSTICE has validated the output product. We are most
appreciative of their timely response.
- Testing HRDI V4.4 and WINDII V5.11A on the Rehost System.
- A full day level zero transmission was received from the LZP test
system. These files will be compared to the production
files received for this same time period.
DAAC DATA:
- Data created in the month of January 2000 has been loaded into the DAAC
Transfer Tables for HALOE, NMC, PEM, and UKMO.
***********
FOT
***********
REPORTING PERIOD
This report covers 11 March, 2000 (Orbit 46443, GMT Day 071) through 17 March
2000 (Orbit 46548, GMT Day 077).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 073/0107 46474 AXIS Off
WINDII 076/0914 46524 On
The beta angle ranged between the angles of 47.9 degrees to 11.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+2.0 degrees on 21 March 2000 (DOY 081).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 07 March 2000 and 17 March 2000 are provided in Appendix D.
The clock error ranged from 11.1 to -12.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 08
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 20
% AXIS = 01
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/021 - 2287.496849
- Transponder B: 00/060 - 2287.499207
GROUND SYSTEM OPERATION
Three (3) operational support problem occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/076 show the observatory data loss to be 566 hours, 21
minutes, 30.937 seconds (an increase of 36 minutes 58.336 seconds since the
last report period).
The increase of 36 minutes 58.336 seconds occurred over three (3) days:
- 13 minutes 34.736 seconds on Day 074. Reason unknown.
- 12 minutes 54.472 seconds on Day 075. Reason unknown.
- 10 minutes 29.128 seconds on Day 076. Reason unknown.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT completed the coding for the three TMONs identified for OBC safehold
Failure, Detection, and Correction. Since this was the first time for this FOT
to use the TMON utility programs, some problems occurred during compiling.
Some of the telemetry mnemonics identified in the TMON algorithms could not be
found by the software. Dick Quinn of Lockheed Martin came by and identified
the problem areas which enabled the FOT to finish the coding. Dick is a lead
systems engineer on TERRA and was the lead systems engineer for UARS and
Landsat 4/5.
Dick's time was limited so the FOT was only able to briefly discuss the
safehold status on UARS with him. He provided some ideas for OBC halt that are
being combined with some of our own. Depending on some additional analysis and
simulator testing, the FOT hopes to deliver a complete safehold package, in the
near future.
The remaining work is as follows:
- Code RTS for use in the UARS Simulator.
- Meet with simulator support and update test scenarios.
- Test TMONs and RTSs for coding errors and system response.
- Develop OBC halt safehold protection plan.
- Submit for Peer Review.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for March 21, 2000 (DOY 081). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 March 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46447 071/0606 00-100 NBTR NBTR Management
46454 071/1725 00-100 NBTR NBTR Management
46455 071/1909 00-101 NBTR NBTR Management
46455 071/1948 00-101 OBC Clk Adj fr 0.00 to 20.6 msec/day
46462 072/0608 00-101 NBTR NBTR Management
46469 072/1728 00-101 NBTR NBTR Management
46469 072/1809 00-103 NBTR NBTR Management
46474 073/0107 00-104 AXIS Off
46477 073/0610 00-103 NBTR NBTR Management
46477 073/0610 00-105 OBC Clk Adj fr 20.6 to 0.00 msec/day
46484 073/1730 00-103 NBTR NBTR Management
46484 073/1812 00-106 NBTR NBTR Management
46492 074/0612 00-106 NBTR NBTR Management
46499 074/1734 00-106 NBTR NBTR Management
46500 074/1955 00-107 NBTR NBTR Management
46507 075/0615 00-107 NBTR NBTR Management
46514 075/1737 00-107 NBTR NBTR Management
46514 075/1816 00-108 NBTR NBTR Management
46522 076/0617 00-108 NBTR NBTR Management
46524 076/0914 00-110 WINDII Off
46529 076/1741 00-109 NBTR NBTR Management
46536 077/0543 00-109 NBTR NBTR Management
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 March 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 46450
PEA1IMONE Y-HI AXIS Input Current Monitor 46469
HRPTSPBIMON R-LO PTS Pulse Bus Current Monitor 46485
HRPSEPBIMON R-LO PSE Pulse Bus Current Monitor 46485
WIQBVOLTAGE Y-LO WINDII Quite Bus Voltage 46
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 46
WIQBVOLTAGE Y-LO WINDII Quite Bus Voltage 46511
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 46511
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None to report this period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 March 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46455 071/19:48 (11 Mar 00) from 0.00 to 20.6 msec/day
46477 073/06:10 (13 Mar 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 March 2000
07 Mar 2000 - GMT Day 067 Beta = 47.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------- ------- ------
1 -50.4/-72.8 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +22.4/-22.4 +5.47/+4.51 +27.2/-15.2 24.8 1.042 21.3
3 +44.8/-5.6 +4.51/+3.23 +26.8/-15.6 24.8 1.040 21.0
17 Mar 2000 - GMT Day 077 Beta = 11.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -44.8/-67.2 -3.05/-4.31 +0.0/+0.0 16.0 0.000 0.0
2 +22.4/-5.6 +5.79/+5.15 +30.0/-12.8 24.2 1.030 22.0
3 +50.4/+11.2 +4.51/+3.87 +28.8/-12.8 24.1 1.027 21.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 March 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-005 FOT 46447 071/0606 14170 Generic Late Acq.
00-005 FOT 46461 072/0428 14170 Generic Late Acq.
00-005 FOT 46514 075/1737 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No closure information received
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-005 FOT see list March 00 14170 Generic Late Acq.
SEVNTFW00/11